Consider a two-way analysis of variance experiment with treatment factors A and B. The results are summarized below.

Source of Variation df SS
Factor A 4 86
Factor B 5 75
Interaction 20 75
Error 90 300
Total 119 536

What are the levels of factor A and factor B?
Answer #1

Ans:

df(factor A)=4

So,number of levels for factor A=4+1=5

df(factor B)=5

So,number of levels for factor B=5+1=6